OpenStudy (anonymous):

r^2-9 factor medals will be awarded

OpenStudy (anonymous):

Alright, you're going to walk through this one with me.

OpenStudy (anonymous):

1/r^7

OpenStudy (anonymous):

Notice that r^2 - 9 is the same thing as r^2 + 0 r -9

OpenStudy (anonymous):

You're looking for two numbers, a and b, that multiply to get -9 and add up to get 0.

OpenStudy (anonymous):

ok...I am tracking...just a little.

OpenStudy (anonymous):

Tracking?

OpenStudy (anonymous):

What are two numbers that multiply together to get -9?

OpenStudy (anonymous):

3 and -3

OpenStudy (anonymous):

Awesome. Do they add up to 0?

OpenStudy (anonymous):

um I think so...that part confuses me a little

OpenStudy (anonymous):

3+(-3) =

OpenStudy (anonymous):

It does equal zero :) . As long as you have two numbers that multiply together to get -9 and add up to get 0, you have the right numbers. We just fill in (x-3)(x+3) and we're done!

OpenStudy (anonymous):

ok...I am still a little confused but a little less than before. Thank you. I will still be posting problems though, they are all the ones I had problems with.

OpenStudy (anonymous):

Alrighty. We can walk through another one. After a couple more, I think you'll have the hang of it. One problem won't be enough, but a couple more might.

OpenStudy (anonymous):

ok thank you...

OpenStudy (anonymous):

Lol, couldn't tell if that was sarcastic or not. The help is there if you want it, but I won't shove it down your throat if you don't.

myininaya (myininaya):

r^2+0*r-9 find two factors of -9 that multiply to be -9 and add up to be 0 -3(3)=-9 and -3+3=0 r^2-3r+3r-9 factor by grouping r(r-3)+3(r-3) (r-3)(r+3)

OpenStudy (anonymous):

Good work guys. I think this one is taught (or learned) better by 'difference of two squares.'